OpenStudy (anonymous):

do anyone know this? f(x) = −3x − 3 and g(x) = 2x + 1, find f over g(−2).

OpenStudy (hba):

g(-2)=2(-2)+1=-4+1=-3

OpenStudy (anonymous):

thanks

OpenStudy (hba):

We have to find f(x)/g(-2)=-3x-3/-3=-3(x+1)/-3=x+1